UT Dec 13 2021 - Split Night DEC/LBTI



Night Info:

  • Sunset 17:28
  • Twlight Ends 18:50
  • Night Center 00:21
  • Moon Set 02:24
  • Twlight Begins 05:51
  • Sunrise 07:14
  • Starting Config LBTI
  • Handover from LBTI to LUCI-1 AO/Eng 20:31hr
  • Handover from LUCI-1 AO/Eng 00:16 (split 10min reconfig between AO/LBTI)

Night Plan

  • LUCI-1 AO Engineering 20:31-00:16hr
    • Bin 3 and BIn 4 stars
    • ESM Bin 2 and Bin 4
  • Poor Weather Backup
    • LUCI-Bino Local Mergers
    • MODS-Bino Local Mergers if LFBG also offline
  • If LFBG offline and conditons are good
    • LBC TMS
      • TMS Setup
        • Service_Observing/2021B_D_E/TMS_13Dec
        • ssh telescope@robs
        • cd /lbt/lbto/supportscripts/MetrologySupport
        • Check if reference file is copied over (J. Hill or O. Kuhn will verify)
        • start ./tms_loop.pl to get mirrors into position
        • CTRL^C the TMS
        • Go to co-point field, run dofpia
        • AFTER dofpia is completed then run TMS
          • First Set reference: ./lbc_set_tms_offset.py -s 2
            • this will take a while before returning prompt
          • In MetrologySupport directory execute ./tms_loop.pl
        • co-point with lbcrangebal, then go to field of choice
        • Start running scripts
        • CTRL-C before running DOFPIA again - if run DOFPIA and TMS simultaneously, the two will "fight" - TMS will backout anything DOFPIA does!

Staff

  • OSA: J. Williams
  • ISp: Mtn
  • Mtn Manager: M.Wagner
  • ISA/SO: B. Rothberg, R.T.Gatto

Night Log

  • UT 00:23 - powering up LBCs on OBS5 as LBTO (will use this to test)
  • UT 00:28 - taking test Biases - OK
  • UT 03:31 - reconfiguring for LBCs
  • UT 03:46 - TMS lasers on
  • UT 03:51 - starting TMS loop
  • UT 03:52 - sending preset to co-pointing field for collimation
  • UT 03:54 - CTRL C'd the tms loop - starting dohybrid
  • UT 04:01 - DIMM ~1.4" along line of sight
  • UT 04:04 - collimation done, running co-pointing
  • offsets
lbcrangebal_1.jpg
  • UT 04:09 - no issue with lbcrangebal (although offsets were large) - H.Choi is working on getting TMS re-set up - there was an error before - will wait for his all clear before setting reference
  • UT 04:14 - H.Choi rebooting the system - need another 3-4min

OB First Round

  • UT 04:18 - setting reference
  • UT 04:19 - starting TMS right only, starting the OB
  • UT 04:27 - FWHM~1.75" on LBC images
  • UT 04:30 - script has other filters (shows 450 images per side) - first set of 25 are g and r-sloan, taking those and will stop OB and run DOFPIA after that (O.Kuhn updating script).
  • UT 04:32 - DIMM ~2"
  • UT 04:40 - LBCs showing ~1.24" DIMM ~ 1.3"
  • UT 04:47 - LBCs~1.19" (for a while LBC-Blue was doing better), DIMM ~1.3"
  • UT 04:51 - seeing ~1.12" on Blue, red is close but still a bit larger FWHM
  • UT 04:59 - blue ~0.92" red is now 0.87" or thereabouts, and red now has slightly smaller FWHMs than blue
  • UT 05:04 - finishing OB - 25 & 24

OB 2nd Round

  • UT 05:04 - running DOFPIA blue only
  • UT 05:08 - starting OB V2
  • UT 05:11 - LBC Red PSFs elongatede, ellpiticity ~0.15 #50852 LBC-R
  • UT 05:13 - #51036 - LBC-R rounder FWHM~1.3", DIMM ~1.4". LBC-B FWHM ~1.2"
  • UT 05:18- LBC-B ~1.08", LBC-R ~1.17"
  • UT 05:19 - LBC-B ~1.03", LBC-R ~ 1.15", DIMM along line of sight ~1.4"
  • UT 05:24 - humidity ~86%
  • #52415- fuzzy trail on LBC-R, don't see it on LBC-B
  • UT 05:33 - LBC images FWHM ~1.1"-1.17" - LBC-B still slightly smaller FWHMs, Humidity 89%
  • UT 05:42 - seeing has gotten better, DIMM 1.0-1.1"
  • UT 05:46 - LBC-B FWHM~1.1", LBC-R FWHM~1.3" DIMM~1.1-1.3", LBC-B PSFs a bt elongated, LBC-R definitely rounder
  • #54624 LBC-R - eleongated PSFs

OB Round 3

  • UT 05:53 - running DOFPIA blue-only
  • UT 05:58 - sending OB
  • UT 05:59 - humidity 87%
  • #55923 LBC-R jumped
  • #60427 LBC-R PSFs look elongated
  • UT 06:11 - seeing on LBC images <1", B-0.7", R-0.78" DIMM~1.1"
  • #61823 LBC-R PSFs elongated
  • #62333 LBC-R PSFs elongated (looks like a jump)
  • #63045 LBC-R PSFs elongated
  • UT 06:37 - humidity dropped to 77%
  • UT 06:34 - LBC-B FWHM~0.63"

OB Round 4

  • UT 06:44 - running DOFPIA, /blueonly
  • UT 06:47 - stariting OB
  • #64831 - no jumps but PSFs do look elongated
  • UT 07:06 - stopped OB, running DOFPIA both sides
  • UT 07:10 - reconfiguring for LBTI
  • UT 07:11 - powering off LBCs


-- BarryRothberg - 12 Dec 2021
Topic revision: r4 - 13 Dec 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback